OpenStudy (anonymous):

8 questions medal and follow

OpenStudy (anonymous):

Question 2 (Essay Worth 4 points) The Right Move home-moving company has a variety of cardboard packing boxes available for use. The packing boxes shown here are similar figures. What is the volume of the larger box? Show your work and explain how you arrived at your answer by applying the scale factor rule of volume. Two rectangular moving boxes are shown side-by-side. The first one has length of two feet, width of zero point five feet, and height of one foot. The second one has length of six feet. Question 3 (Multiple Choice Worth 2 points) The ratio of corresponding sides of similar plane figures is twenty-eight inches divided by seven inches. If the area of the smaller figure is fifteen square inches, what is the area of the larger figure? 36 sq ft. 60 sq ft 120 sq ft. 240 sq ft. Question 4 (Multiple Choice Worth 2 points) After using a new rectangular cloth napkin, Mister Weston puts it in his washing machine, then in the dryer. When it comes out, its length and width are three-fourths their original sizes. How has the area of the napkin changed? Two napkins are shown. The first one is the original napkin. The second one is the napkin that has shrunk, which appears to be three-fourths the size of the original napkin. It has been reduced to three-fourths its original area. It has been reduced to three-eighths its original area. It has been reduced to nine-sixteenths its original area. It has been reduced to twenty-seven sixty-fourths its original area. Question 5 (Multiple Choice Worth 2 points) Trapezoid P Q R S and trapezoid W X Y Z shown below are similar. What is the perimeter of trapezoid W X Y Z? Trapezoid P Q R S and trapezoid W X Y Z are placed side-by-side. Side P Q is nine meters. Side Q R is twelve meters. Side R S is nine meters. Side S P is six meters. Side X Y is eight meters. 6 meters twenty meters twenty-four meters thirty-six meters Question 6 (Multiple Choice Worth 2 points) The area of triangle A B C is twelvesquare centimeters. Each of the dimensions of triangle D E F is 3 times the dimension of triangle A B C. What is the area of triangle D E F? Triangle A B C and triangle D E F are placed side-by-side. They are both similar triangles. Triangle A B C is smaller than triangle D E F. 4 square centimeters thirty-six square centimeters seventy-two square centimeters one hundred eight square centimeters Question 7 (Multiple Choice Worth 2 points) If the side length of the cube shown here is changed by a scale factor of 5, what will be the volume of the new cube? A cube with side length of four inches is shown. three hundred twenty cubic inches four hundred cubic inches nine hundred sixty cubic inches eight thousand cubic inches Question 8 (Multiple Choice Worth 2 points) The volume of a cylinder increased by 8 times the original volume. How many times greater are each of the dimensions of that new cylinder? 2 3 4 8

OpenStudy (anonymous):

@hippo4

OpenStudy (anonymous):

@Warriorz13 @YanaSidlinskiy @whpalmer4 @adrynicoleb @inkyvoyd @bbcream14 @yo_its_taylorrrrrrr

OpenStudy (anonymous):

@erocdrah72 @TheOcean @TheRaggedyDoctor @Ufukedis @karategirl2002 @ilovebmth1234 @minnie♥mouse

OpenStudy (anonymous):

@MelissaHolmes

OpenStudy (anonymous):

im not good with math sorry kiddo... :/

OpenStudy (yanasidlinskiy):

This is for number 1. Please post one question at a time, than tacking them all at once Well, by what you said we know they are similar, right? If they are similar then the ratio of one side of one box to the corresponding side of the other box is the same for all sides of the box. We know one side of the bigger box is 6, and the corresponding (the side that is the same side on the smaller box) is 2. So to get from the side of the smaller box (2) to the side of the bigger box (6) you multiply by 3. So, since this 3 is the scale factor. Since it is always the same do that same thing for all sides of small box to find the other sides of the bigger box. So the three sides of the bigger box will be: 2*3=6 .5*3=1.5 1*3=3 The volume of the small box is length*width*height, so that is 2*.5*3=3 The volume of the bigger box is length*width*height, so that is 6*1.5*3=27

OpenStudy (yanasidlinskiy):

Do you understand?